OpenStudy (anonymous):

trig... same stuff I have been doing... can some one check my work? I dont know where I went wrong....

OpenStudy (anonymous):

where is the question?

OpenStudy (freckles):

you want 4 roots so first realize sin and cos have period 2pi \[z=81(\cos(\frac{4\pi}{3}+2 npi)+i \sin(\frac{4\pi}{3}+2 n \pi))\] now you want the 4th root of this so you will do the whole n=0,1,2,3 thing where \[z^\frac{1}{4}=81^\frac{1}{4}(\cos(\frac{1}{4}[\frac{4\pi}{3}+2n \pi])+i \sin[ \frac{1}{4}[\frac{4\pi}{3}+2n \pi]))\]

OpenStudy (freckles):

let me look at your work

OpenStudy (freckles):

so n=0 we have \[3(\cos(\frac{\pi}{3})+i \sin(\frac{\pi}{3}))\] you did good there for n=1 we have \[3(\cos(\frac{5\pi}{6})+i \sin(\frac{5\pi}{6}))\] that one also looks good now for n=2 we have that looks good I think your n=3 is off just a hair

OpenStudy (freckles):

you change 3 to 2?

OpenStudy (freckles):

also \[3(\frac{\sqrt{3}}{2}-\frac{1}{2}i)=\frac{3 \sqrt{3}}{2}-\frac{3}{2}i\] is this what you have for n=3?

OpenStudy (freckles):

you did really good I just think you got tired of all that writing :p

OpenStudy (anonymous):

ok let me see if it works..

OpenStudy (anonymous):

nope something is wrong.. because the software keeps marking it wrong

OpenStudy (freckles):

everything on your paper is good up until that last line you have

OpenStudy (freckles):

like you corrected just the last line right?

OpenStudy (anonymous):

oh wait i think i put it in wrong.

OpenStudy (anonymous):

i wasnt putting in the answer in the right order! silly silly mistakes